# Community detection in the hypergraph stochastic block model and reconstruction on hypertrees

### v1

@@ Abstract
We study the weak recovery problem on the $r$-uniform hypergraph stochastic block model ($r$-HSBM) with two balanced communities. In this model, $n$ vertices are randomly divided into two communities, and size-$r$ hyperedges are added randomly depending on whether all vertices in the hyperedge are in the same community. The goal of the weak recovery problem is to recover a non-trivial fraction of the communities given the hypergraph. Previously, [PZ21] established that weak recovery is always possible above a natural threshold called the Kesten-Stigum (KS) threshold. [GP23b] proved that the KS threshold is tight if $r\leq 4$ or the expected degree $d$ is small. It remained open whether the KS threshold is tight for $r\geq 5$ and large $d$.
In this paper we determine the tightness of the KS threshold for any fixed $r$ and large $d$. We prove that for $r\leq 6$ and $d$ large enough, the KS threshold is tight. This shows that there is no information-computation gap in this regime. This partially confirms a conjecture of [ACKZ15]. For $r\geq 7$, we prove that for $d$ large enough, the KS threshold is not tight, providing more evidence supporting the existence of an information-computation gap in this regime.
Furthermore, we establish asymptotic bounds on the weak recovery threshold for fixed $r$ and large $d$. We also obtain a number of results regarding the closely-related broadcasting on hypertrees (BOHT) model, including the asymptotics of the reconstruction threshold for $r\geq 7$ and impossibility of robust reconstruction at criticality.
@@ 1.1 Hypergraph stochastic block model
The stochastic block model (SBM) is a random graph model with communities structures. It is the canonical model for studying community detection, and has applications in statistics, machine learning, and network science. The theoretical study of community detection in the SBM has received a lot of attention in the past decade since [DKMZ11]; see [Abb17] for a survey.
The hypergraph stochastic block model (HSBM) is a generalization of the SBM to hypergraphs. In this paper, we focus on the $r$-uniform HSBM with two symmetric communities, defined as follows. Let $n\in\mathbb{Z}_{\geq 0}$ denote the number of vertices and $V=[n]$ be the set of vertices [^1] We use the notation $[n]=\{1,\ldots,n\}$.. Let $r\in\mathbb{Z}_{\geq 2}$ be the hyperedge size and $a,b\in\mathbb{R}_{\geq 0}$ be two parameters [^2] Several previous works (e.g., [PZ21, ZSTG22, GP23b]) assume that $a\geq b$, but we do not make this assumption.. Generate a random label $X_{u}\in\{\pm\}$ for all vertices $u\in V$ i.i.d. $\sim\operatorname{\mathrm{Unif}}(\{\pm\})$. For every subset $S\subseteq V$ of size $r$, the hyperedge $S$ is added with probability $\frac{a}{\binom{n}{r-1}}$ if all vertices in $S$ have the same label; otherwise add the hyperedge with probability $\frac{b}{\binom{n}{r-1}}$. The model outputs $(X,G)$, where $G$ is the resulting hypergraph. We denote the model as $\operatorname{\mathrm{HSBM}}(n,r,a,b)$.
@@
where
$d_{H}(X,Y)=\min_{s\in\{\pm\}}\sum_{u\in V}\mathbbm{1}\{X_{i}\neq sY_{i}\}.$ (3)
For $r=2$, the model $\operatorname{\mathrm{HSBM}}(n,2,a,b)$ becomes SBM with two symmetric communities, where the weak recovery threshold was established by [Mas14, MNS15, MNS18]: weak recovery is possible if and only if $d\lambda^{2}>1$. The threshold $d\lambda^{2}=1$ is called the Kesten-Stigum (KS) threshold and is originated from the study of multitype Galton-Watson processes ([KS66]). For the HSBM, the natural generalization of the KS threshold is $(r-1)d\lambda^{2}=1$. [ACKZ15] conjectured that a phase transition for weak recovery occurs at the KS threshold. [PZ21] proved the positive part of the conjecture, that is, when $(r-1)d\lambda^{2}>1$, weak recovery is possible and can be achieved using an efficient (polynomial time) algorithm. On the negative part, [GP23b] proved that when $r=3,4$ or $\lambda\geq\frac{1}{5}$, if $(r-1)d\lambda^{2}\leq 1$, then weak recovery is impossible. They also provided evidence (via studying the reconstruction on hypertrees problem, see Section 1.2) that for $r\geq 7$ and $d$ large enough, weak recovery might be information-theoretically possible slightly below the KS threshold. Nevertheless, the tightness of the KS threshold for $r\geq 5$ and large $d$ remained open before the current work.
Our main contribution is to determine the tightness of the KS threshold for every $r$ and large $d$. We show that the behavior is different for $r\leq 6$ and $r\geq 7$: for $r\leq 6$ and large $d$, the KS threshold is tight; for $r\geq 7$ and large $d$, the KS threshold is not tight.
@@
For $r\geq 7$, we prove asymptotic lower and upper bounds on the HSBM weak recovery threshold.
@@
Note that $\frac{2r\log 2}{2^{r-1}-1}\leq\frac{1}{r-1}$ for $r\geq 7$. Therefore for $r\geq 7$, the KS threshold is not tight.
Corollary 3 (KS is not tight for $r\geq 7$ and large degree, HSBM).
Work in the setting of Theorem 2. For any $r\geq 7$, there exists $d,\lambda$ satisfying $(r-1)d\lambda^{2}<1$ such that weak recovery is possible.
Our current knowledge about the tightness of the KS threshold for HSBM weak recovery is summarized in Table 1. Our contribution is the “large $d$” column (except for the $r=3,4$, $\lambda>0$ case which was established in [GP23b]).
@@
$r=3,4$, $\lambda>0$ tight (tight) (tight)
@@
$r=5,6$, $\lambda<0$ unknown unknown tight
$r\geq 7$, $\lambda>0$ (not tight) tight not tight
$r\geq 7$, $\lambda<0$ (not tight) unknown not tight
Table 1: Tightness of KS threshold for HSBM weak recovery. An entry in parentheses means it is implied by another entry in the same row. References: Possibility above KS: [PZ21] ($\lambda>0$), [SZ22] (general HSBM). Tightness for $r=3,4$, $\lambda>0$, and $r\geq 5$, $\lambda>0$, small $d$: [GP23b]. Tightness for $r\leq 6$, large $d$: Theorem 1. Non-tightness for $r\geq 7$, large $d$: Corollary 3.
Proof of Theorem 1 and Theorem 2(ii) are based on the reconstruction on hypertrees problem, which we will discuss in Section 1.2.
Theorem 2(i) follows from a non-asymptotic weak recovery upper bound, which we now describe.
For $x,y\in[0,1]$, define the binary Kullback-Leibler (KL) divergence function $d_{\operatorname{\mathrm{KL}}}(\cdot\|\cdot)$ as [^3] Throughout the paper, $\log$ denotes natural logarithm.
We obtain the following non-asymptotic upper bound for weak recovery.
Theorem 4 (Non-asymptotic HSBM weak recovery upper bound).
then weak recovery is possible.
Theorem 2(i) follows from Theorem 4 by taking the $\lambda\to 0$ limit.
Proof of Theorem 4 uses a simple estimator which returns a balanced partition with high in-community hyperedge density. We show that when Eq. (5) is satisfied, any balanced partition that achieves high in-community hyperedge density is correlated with the true partition. This method was previously used in [BMNN16] in the case of $q$-community SBM.
@@ 1.2 Reconstruction on hypertrees
Theorem 1 and Theorem 2(ii) are proved by studying the reconstruction problem for the broadcasting on hypertrees (BOHT) model, defined as follows. Let $r\in\mathbb{Z}_{\geq 2}$, be the hyperedge size, $\lambda\in\left[-\frac{1}{2^{r-1}-1},1\right]$ be a strength parameter, $D$ be a distribution on $\mathbb{Z}_{\geq 0}$, which is usually taken to be the point distribution at $d\in\mathbb{Z}_{\geq 0}$ or the Poisson distrbution $\operatorname{\mathrm{Pois}}(d)$ with expectation $d\in\mathbb{R}_{\geq 0}$. Let $T$ be the $r$-uniform Galton-Watson linear [^4] Linear means the intersection of two different hyperedges has size at most one. hypertree, where every vertex independently has $b\sim D$ downward hyperedges. The BOHT model assigns a $\pm$ label to every vertex according to the following process. The root $\rho$ has label $\sigma_{\rho}$ generated from the uniform distribution $\pi=\operatorname{\mathrm{Unif}}(\{\pm\})$. Let $B_{r,\lambda}:\{\pm\}\to\{\pm\}^{r-1}$ be the channel
@@
Suppose we have generated label $\sigma_{u}$ of a vertex $u$. Then for every downward hyperedge $\{u,v_{1},\ldots,v_{r-1}\}$, children labels $\sigma_{v_{1}},\ldots,\sigma_{v_{r-1}}$ are generated according to $B_{r,\lambda}$, that is, $(\sigma_{v_{1}},\ldots,\sigma_{v_{r-1}})\sim B_{r,\lambda}(\cdot|\sigma_{u})$. In this way every vertex has received a label. The model outputs $(T,\sigma)$ and we denote the model as $\operatorname{\mathrm{BOHT}}(r,\lambda,D)$. When $D$ is a point distrbution at $d\in\mathbb{Z}_{\geq 0}$ we denote the model as $\operatorname{\mathrm{BOHT}}(r,\lambda,d)$.
The reconstruction problem is a very important problem for the BOHT model. Let $(T,\sigma)\sim\operatorname{\mathrm{BOHT}}(r,\lambda,D)$. For $k\in\mathbb{Z}_{\geq 0}$, let $T_{k}$ denote the tree structure up to the $k$-th level (the root being at the 00-th level), and $L_{k}$ denote the set of vertices at the $k$-th level. We say reconstruction is possible if
@@
is positive, and impossible otherwise.
The reconstruction problem is useful for the study of HSBM weak recovery. Given a model $\operatorname{\mathrm{HSBM}}(n,r,a,b)$. Let $d,\lambda$ be defined as in Eq. (1). Consider the model $\operatorname{\mathrm{BOHT}}(r,\lambda,\operatorname{\mathrm{Pois}}(d))$, which we call the BOHT model corresponding to $\operatorname{\mathrm{HSBM}}(n,r,a,b)$. If reconstruction is impossible for $\operatorname{\mathrm{BOHT}}(r,\lambda,\operatorname{\mathrm{Pois}}(d))$, then weak recovery is impossible for $\operatorname{\mathrm{HSBM}}(n,r,a,b)$. This relationship was first established by [MNS15, MNS18] for the $r=2$ case, and their proof can be easily generalized. See [Gu23, Theorem 5.15] for a proof for general HSBMs. We note that the converse is not true in general: for the $q$-community SBM, there exist parameters where weak recovery is impossible for the SBM but reconstruction is possible for the corresponding broadcasting on trees (BOT) model.
In the $r=2$ case, the model $\operatorname{\mathrm{BOHT}}(2,\lambda,D)$ is called the Ising model, where the reconstruction threshold was established by [BRZ95, EKPS00]: reconstruction is possible if and only if $d\lambda^{2}>1$. The threshold $d\lambda^{2}=1$ is also called the Kesten-Stigum (KS) threshold. For $r\geq 3$, the natural generalization of the KS threshold is $(r-1)d\lambda^{2}=1$. In general, it is known that reconstruction is possible whenever $(r-1)d\lambda^{2}>1$, and the tightness of the KS threshold is an interesting question. [GP23b] proved that for reconstruction for $\operatorname{\mathrm{BOHT}}(r,\lambda,D)$ is impossible for $(r-1)d\lambda^{2}\leq 1$ if $r\leq 4$, $\lambda>0$, or $\lambda\geq\frac{1}{5}$. Furthermore, they prove that the KS threshold is not tight for $r\geq 7$ and $d$ large enough. The remaining cases were left open.
Our main contribution for the BOHT model is tightness of the KS threshold for $r\leq 6$ and $d$ large enough.
@@
Not coincidentally, Table 1 is also our current knowledge about tightness of the KS threshold for the reconstruction problem on $\operatorname{\mathrm{BOHT}}(r,\lambda,D)$ (for $D=d$ or $D=\operatorname{\mathrm{Pois}}(d)$). Possibility of reconstruction above KS is folklore by using a counting estimator. Tightness for $r=3,4$, $\lambda>0$, and $r\geq 5$, $\lambda>0$, small $d$, and non-tightness for $r\geq 7$, large $d$ are by [GP23b]. Tightness for $r\leq 6$, large $d$ (except for $r=3,4$, $\lambda>0$) is by Theorem 5.
Theorem 1 directly follows from Theorem 5 by using the relationship between HSBM weak recovery and BOHT reconstruction.
@@
(See Section 2.1 for relevant information theory background.) In other words, $M_{k}=\operatorname{\mathrm{BP}}^{k}(\operatorname{\mathrm{Id}})$, where $\operatorname{\mathrm{Id}}:\{\pm\}\to\{\pm\}$ denotes the identity channel. Then non-reconstruction is equivalent to
@@
In the first part (large degree asymptotics), we prove that for $r\leq 6$ and any $\epsilon>0$, there exists $d_{0}=d_{0}(r,\epsilon)$ such that for $d\geq d_{0}$ and $(r-1)d\lambda^{2}\leq 1$, we have
$\lim_{k\to\infty}C_{\chi^{2}}(\operatorname{\mathrm{BP}}^{k}(\operatorname{\mathrm{Id}}))\leq\epsilon,$ (13)
where $C_{\chi^{2}}$ denotes $\chi^{2}$-capacity. The idea is that $C_{\chi^{2}}(P^{\star d})$ for a BMS channel $P$ is the expectation of a function involving $d$ i.i.d. variables. So when $d$ is large, the result can be approximated using the expectation of a function involving a Gaussian variable. For the BOHT model, [GP23b] established the Gaussian approximation. The resulting function is a Gaussian integral (Eq. (23)), and we prove the numerical observation that it has no fixed points on $(0,1]$.
The second part (robust reconstruction) is more difficult. We prove that for any $r\geq 2$, there exists $\epsilon=\epsilon(r)>0$ such that if $(r-1)d\lambda^{2}\leq 1$, then for any BMS channel $P$ with $C_{\chi^{2}}(P)\leq\epsilon$, we have
@@
There are several difficulties in the this approach. The first difficulty is that $C_{\chi^{2}}(\operatorname{\mathrm{BP}}(P))$ is not a function of $C_{\chi^{2}}(P)$. Any BMS $P$ is described by a random variable $\theta_{P}$ on $[-1,1]$ (see Section 2.1). The $\chi^{2}$-information $C_{\chi^{2}}(P)$ satisfies
@@
Because $\|\theta_{P}\|\leq 1$ almost surely for any $P$, we have $y_{k}\leq x_{k}$. We first perform a crude expansion of $x_{k+1}$ and $y_{k+1}$, which roughly speaking allows us to conclude that $y_{k}=O(x_{k}^{2})$. In this way we know that $y_{k}$ does not affect the first two orders if we perform a finer expansion of $x_{k+1}$. The finer expansion will give us Eq. (15).
The second difficulty is that, unlike previous works which work with trees, we work with hypertrees. As we noted, $C_{\chi^{2}}(\operatorname{\mathrm{BP}}(P))$ is the expectation of a function with $(r-1)d$ variables, while previous works have $d$ variables. When performing the expansion, there will be a large number of terms, even after considering symmetry. Our solution is to perform the expansion in two steps. Let $M^{\prime}_{k+1}$ be the channel $M_{k}^{\times(r-1)}\circ B_{r,\lambda}$. Define
@@
Then we can decompose the expansion step $(x_{k},y_{k})\to(x_{k+1},y_{k+1})$ into two smaller steps $(x_{k},y_{k})\to(x^{\prime}_{k+1},y^{\prime}_{k+1})$ and $(x^{\prime}_{k+1},y^{\prime}_{k+1})\to(x_{k+1},y_{k+1})$. The fine expansions we get for $x$ and $x^{\prime}$ are as follows.

### v2

@@ Abstract
We study the weak recovery problem on the $r$-uniform hypergraph stochastic block model ($r$-HSBM) with two balanced communities. In this model, $n$ vertices are randomly divided into two communities, and size-$r$ hyperedges are added randomly depending on whether all vertices in the hyperedge are in the same community. The goal of weak recovery is to recover a non-trivial fraction of the communities given the hypergraph. [PZ21, SZ22] established that weak recovery is always possible above a natural threshold called the Kesten-Stigum (KS) threshold. For assortative models (i.e., monochromatic hyperedges are preferred), [GP23b] proved that the KS threshold is tight if $r\leq 4$ or the expected degree $d$ is small. For other cases, the tightness of the KS threshold remained open.
In this paper we determine the tightness of the KS threshold for a wide range of parameters. We prove that for $r\leq 6$ and $d$ large enough, the KS threshold is tight. This shows that there is no information-computation gap in this regime and partially confirms a conjecture of [ACKZ15]. On the other hand, we show that for $r\geq 5$, there exist parameters for which the KS threshold is not tight. In particular, for $r\geq 7$, the KS threshold is not tight if the model is disassortative (i.e., polychromatic hyperedges are preferred) or $d$ is large enough. This provides more evidence supporting the existence of an information-computation gap in these cases.
Furthermore, we establish asymptotic bounds on the weak recovery threshold for fixed $r$ and large $d$. We also obtain a number of results regarding the broadcasting on hypertrees (BOHT) model, including the asymptotics of the reconstruction threshold for $r\geq 7$ and impossibility of robust reconstruction at criticality.
@@ 1.1 Hypergraph stochastic block model
The stochastic block model (SBM) is a random graph model with community structures. It is the canonical model for studying community detection, and has applications in statistics, machine learning, and network science. The theoretical study of community detection in the SBM has received a lot of attention in the past decade since [DKMZ11]; see [Abb17] for a survey.
The hypergraph stochastic block model (HSBM) is a generalization of the SBM to hypergraphs. In this paper, we study the $r$-uniform HSBM with two symmetric communities, defined as follows. Let $n\in\mathbb{Z}_{\geq 0}$ denote the number of vertices and $V=[n]$ be the set of vertices [^1] We use the notation $[n]=\{1,\ldots,n\}$.. Let $r\in\mathbb{Z}_{\geq 2}$ be the hyperedge size and $a,b\in\mathbb{R}_{\geq 0}$ be two parameters [^2] While several previous works (e.g., [PZ21, ZSTG22, GP23b]) assume that $a\geq b$, we do not make this assumption.. Generate a random label $X_{u}\in\{\pm\}$ for all vertices $u\in V$ i.i.d. $\sim\operatorname{\mathrm{Unif}}(\{\pm\})$. For every subset $S\subseteq V$ of size $r$, the hyperedge $S$ is added with probability $\frac{a}{\binom{n}{r-1}}$ if all vertices in $S$ have the same label; otherwise add the hyperedge with probability $\frac{b}{\binom{n}{r-1}}$. The model outputs $(X,G)$, where $G$ is the resulting hypergraph. We denote the model as $\operatorname{\mathrm{HSBM}}(n,r,a,b)$.
@@
$\text{where}\qquad d_{H}(X,Y)=\min_{s\in\{\pm\}}\sum_{u\in V}\mathbbm{1}\{X_{u}\neq sY_{u}\}.$ (3)
For $r=2$, the model $\operatorname{\mathrm{HSBM}}(n,2,a,b)$ becomes SBM with two symmetric communities, where the weak recovery threshold was established by [Mas14, MNS15, MNS18]: weak recovery is possible if and only if $d\lambda^{2}>1$. The threshold $d\lambda^{2}=1$ is called the Kesten-Stigum (KS) threshold and originates from the study of multitype Galton-Watson processes ([KS66]). For the HSBM, the natural generalization of the KS threshold is $(r-1)d\lambda^{2}=1$. [ACKZ15] conjectured that a phase transition for weak recovery occurs at the KS threshold. [PZ21, SZ22] proved the positive part of the conjecture, that is, when $(r-1)d\lambda^{2}>1$, weak recovery is possible and can be achieved using an efficient (polynomial time) algorithm. On the negative part, [GP23b] proved that for assortative models (i.e., $a>b$), when $r=3,4$ or $\lambda\geq\frac{1}{5}$, if $(r-1)d\lambda^{2}\leq 1$, then weak recovery is impossible. They also provided evidence (via studying the reconstruction on hypertrees problem, see Section 1.2) suggesting that for $r\geq 7$ and $d$ large enough, weak recovery might be information-theoretically possible below the KS threshold. Nevertheless, the tightness of the KS threshold for disassortative models (i.e., $a<b$) or assortative models with $r\geq 5$ and large $d$ remained open before the current work.
Our first contribution is to show that the KS threshold is tight for $r\leq 6$ and $d$ large enough.
@@
Our second contribution is to provide an upper bound for HSBM weak recovery. For $x,y\in[0,1]$, define the binary Kullback-Leibler (KL) divergence function $d_{\operatorname{\mathrm{KL}}}(\cdot\|\cdot)$ as [^3] Throughout the paper, $\log$ denotes natural logarithm.
Theorem 2 (HSBM weak recovery upper bound).
then weak recovery is possible.
Comparing Eq. (5) with the KS threshold $(r-1)d\lambda^{2}=1$, we obtain that the KS threshold is not tight for a wide range of parameters.
Corollary 3 (Non-tightness of the KS threshold for $r\geq 5$, HSBM).
Work in the setting of Theorem 2. If $r\geq 5$, then there exists $\lambda_{0}=\lambda_{0}(r)\in\left(-\frac{1}{2^{r-1}-1},1\right]$ such that for any non-zero $\lambda\in\left[-\frac{1}{2^{r-1}-1},\lambda_{0}\right)$, there exists $d$ satisfying $(r-1)d\lambda^{2}<1$ such that weak recovery is possible. Furthermore, for $r\geq 7$ we can take $\lambda_{0}(r)>0$.
In Corollary 3, we can take $\lambda_{0}(r)$ to be the minimum solution in $\left[-\frac{1}{2^{r-1}-1},1\right]$ to the equation
$d_{\operatorname{\mathrm{KL}}}\left(\lambda+\frac{1-\lambda}{2^{r-1}}\left\|\frac{1}{2^{r-1}}\right.\right)=r(r-1)\lambda^{2}\log 2.$ (6)
Table 1 presents numerical values of $\lambda_{0}(r)$ for small $r$.
$r$ $5$ $6$ $7$ $8$ $9$ $10$
$\lambda_{0}$ $-0.06575$ $-0.02154$ $0.00413$ $0.01920$ $0.02807$ $0.03320$
Table 1: Minimum solution to Eq. (6) for $5\leq r\leq 10$, rounded down to the nearest multiple of $10^{-5}$. For non-zero $\lambda\in\left[-\frac{1}{2^{r-1}-1},\lambda_{0}\right)$, the KS threshold is not tight. See Corollary 3 for the complete statement.
Furthermore, we obtain the following asymptotic lower and upper bounds for the HSBM weak recovery threshold.
@@
Our current knowledge about the tightness of the KS threshold for HSBM weak recovery is summarized in Table 2. Our contribution is the “large $d$” column (except for the $r=3,4$, $\lambda>0$ case which was established in [GP23b]) and all non-tightness results.
@@
$r=3,4$, $\lambda>0$ always tight (tight) (tight)
@@
$r=5,6$, $\lambda<0$ unknown not tight tight
$r\geq 7$, $\lambda>0$ unknown tight not tight
$r\geq 7$, $\lambda<0$ never tight (not tight) (not tight)
Table 2: Tightness of the KS threshold for HSBM weak recovery. An entry in parentheses means it is implied by another entry in the same row. References: Possibility above KS: [PZ21] ($\lambda>0$), [SZ22] (general HSBM). Tightness for $r=3,4$, $\lambda>0$, and $r\geq 5$, $\lambda>0$, small $d$: [GP23b]. Tightness for $r\leq 6$, large $d$: Theorem 1. Non-tightness results: Corollary 3. This table is also our current knowledge about tightness of the KS threshold for BOHT reconstruction.
Proofs of Theorem 1 and Theorem 4(ii) are based on the reconstruction on hypertrees problem, which we will discuss in Section 1.2. Proof of Theorem 2 uses a simple estimator which returns a balanced partition with expected in-community hyperedge density. We show that when Eq. (5) is satisfied, any balanced partition that achieves expected in-community hyperedge density is correlated with the true partition. This method was previously used in [BMNN16] in the case of $q$-community SBM. Corollary 3 and Theorem 4(i) follow from Theorem 2.
Another interesting community detection problem for the HSBM is the testing problem, which asks to distinguish an HSBM hypergraph from an Erdős-Rényi hypergraph with the same hyperedge density. [YLFS22] studied the testing problem for a class of symmetric HSBMs (which includes our two-community case), proving that testing (with $o(1)$ error probability) is possible above the KS threshold. They showed that the KS threshold is tight for the two-community three-uniform HSBM, and that the KS threshold is not tight in general. In fact, they proved that when Eq. (5) holds, testing is possible. [^4] We remark that [YLFS22, Theorem 2.6, Part 1, second half] is off by a factor of two because of a computation error. In particular, in the second-to-last display in the proof of Theorem 2.6, the third step performs Taylor expansion to order $\epsilon$. However, because the final result is of order $\epsilon^{2}$, to get the correct coefficient in the final result, one needs to perform Taylor expansion to order $\epsilon^{2}$. Their results do not imply our Theorem 2 because a direct relationship between testing and weak recovery is not known. Nevertheless, this provides evidence suggesting that there might be a close relationship between the two problems.
@@ 1.2 Reconstruction on hypertrees
Theorem 1 and Theorem 4(ii) are proved by studying the reconstruction problem for the broadcasting on hypertrees (BOHT) model, defined as follows. Let $r\in\mathbb{Z}_{\geq 2}$ be the hyperedge size, $\lambda\in\left[-\frac{1}{2^{r-1}-1},1\right]$ be a strength parameter, $D$ be a distribution on $\mathbb{Z}_{\geq 0}$, which is usually taken to be the point distribution at $d\in\mathbb{Z}_{\geq 0}$ or the Poisson distribution $\operatorname{\mathrm{Pois}}(d)$ with expectation $d\in\mathbb{R}_{\geq 0}$. Let $T$ be the $r$-uniform Galton-Watson linear [^5] Linear means the intersection of two different hyperedges has size at most one. hypertree, where every vertex independently has $b\sim D$ downward hyperedges. The BOHT model assigns a $\pm$ label to every vertex according to the following process. The root $\rho$ has label $\sigma_{\rho}$ generated from the uniform distribution $\pi=\operatorname{\mathrm{Unif}}(\{\pm\})$. Let $B_{r,\lambda}:\{\pm\}\to\{\pm\}^{r-1}$ be the channel
@@
Suppose we have generated label $\sigma_{u}$ of a vertex $u$. Then for every downward hyperedge $\{u,v_{1},\ldots,v_{r-1}\}$, children labels $\sigma_{v_{1}},\ldots,\sigma_{v_{r-1}}$ are generated according to $B_{r,\lambda}$, that is, $(\sigma_{v_{1}},\ldots,\sigma_{v_{r-1}})\sim B_{r,\lambda}(\cdot|\sigma_{u})$. In this way every vertex has received a label. The model outputs $(T,\sigma)$ and we denote the model as $\operatorname{\mathrm{BOHT}}(r,\lambda,D)$. When $D$ is a point distribution at $d\in\mathbb{Z}_{\geq 0}$ we denote the model as $\operatorname{\mathrm{BOHT}}(r,\lambda,d)$.
The reconstruction problem is an important problem for the BOHT model. Let $(T,\sigma)\sim\operatorname{\mathrm{BOHT}}(r,\lambda,D)$. For $k\in\mathbb{Z}_{\geq 0}$, let $T_{k}$ denote the tree structure up to the $k$-th level (the root being at the 00-th level), and $L_{k}$ denote the set of vertices at the $k$-th level. We say reconstruction is possible if
@@
is positive, and impossible otherwise. Here $I(\cdot;\cdot)$ denotes mutual information. In other words, reconstruction is possible if and only if we can reconstruct the root label $\sigma_{\rho}$ with $\frac{1}{2}-\Omega(1)$ error probability given the leaf labels $\sigma_{L_{k}}$, as $k\to\infty$.
The reconstruction problem is useful for the study of HSBM weak recovery. Consider the model $\operatorname{\mathrm{HSBM}}(n,r,a,b)$. Let $d,\lambda$ be defined as in Eq. (1). Consider the model $\operatorname{\mathrm{BOHT}}(r,\lambda,\operatorname{\mathrm{Pois}}(d))$, which we call the BOHT model corresponding to $\operatorname{\mathrm{HSBM}}(n,r,a,b)$. If reconstruction is impossible for $\operatorname{\mathrm{BOHT}}(r,\lambda,\operatorname{\mathrm{Pois}}(d))$, then weak recovery is impossible for $\operatorname{\mathrm{HSBM}}(n,r,a,b)$. This relationship was first established by [MNS15, MNS18] for the $r=2$ case, and their proof can be easily generalized. See [Gu23, Theorem 5.15] for a proof for general HSBMs. We note that the converse is not true in general: for the $q$-community SBM, there exist parameters where weak recovery is impossible for the SBM but reconstruction is possible for the corresponding broadcasting on trees (BOT) model.
In the $r=2$ case, the model $\operatorname{\mathrm{BOHT}}(2,\lambda,D)$ is called the Ising model, where the reconstruction threshold was established by [BRZ95, EKPS00]: reconstruction is possible if and only if $d\lambda^{2}>1$. The threshold $d\lambda^{2}=1$ is also called the Kesten-Stigum (KS) threshold. For $r\geq 3$, the natural generalization of the KS threshold is $(r-1)d\lambda^{2}=1$. In general, it is known that reconstruction is possible whenever $(r-1)d\lambda^{2}>1$, and the tightness of the KS threshold is an interesting question. [GP23b] proved that reconstruction for $\operatorname{\mathrm{BOHT}}(r,\lambda,D)$ is impossible for $(r-1)d\lambda^{2}\leq 1$ if $r\leq 4$, $\lambda>0$, or $\lambda\geq\frac{1}{5}$. Furthermore, they proved that the KS threshold is not tight for $r\geq 7$ and $d$ large enough. The remaining cases were left open.
Our main contribution on the BOHT model is tightness of the KS threshold for $r\leq 6$ and large $d$.
@@
Not coincidentally, Table 2 is also our current knowledge about tightness of the KS threshold for the reconstruction problem on $\operatorname{\mathrm{BOHT}}(r,\lambda,D)$ (for $D=d$ or $D=\operatorname{\mathrm{Pois}}(d)$). Possibility of reconstruction above KS is folklore by a counting estimator. Tightness for $r=3,4$, $\lambda>0$, and $r\geq 5$, $\lambda>0$, small $d$, and non-tightness for $r\geq 7$, large $d$ are by [GP23b]. Tightness for $r\leq 6$, large $d$ (except for $r=3,4$, $\lambda>0$) is by Theorem 5. Other non-tightness results are by Corollary 3.
Theorem 1 follows directly from Theorem 5 using the relationship between HSBM weak recovery and BOHT reconstruction ([Gu23, Theorem 5.15]).
@@
(See Section 2.1 for the relevant information theory background.) In other words, $M_{k}=\operatorname{\mathrm{BP}}^{k}(\operatorname{\mathrm{Id}})$, where $\operatorname{\mathrm{Id}}:\{\pm\}\to\{\pm\}$ denotes the identity channel. Then non-reconstruction is equivalent to
@@
In the first part (large degree asymptotics), we prove the following result.
Proposition 6 (Large degree asymptotics).
Fix $r\leq 6$. For any $\epsilon>0$, there exists $d_{0}=d_{0}(r,\epsilon)$ such that for any $d\geq d_{0}$ and $\lambda\in\left[-\frac{1}{2^{r-1}-1},1\right]$ with $(r-1)d\lambda^{2}\leq 1$, we have
$\lim_{k\to\infty}C_{\chi^{2}}(\operatorname{\mathrm{BP}}^{k}(\operatorname{\mathrm{Id}}))<\epsilon,$ (14)
where $C_{\chi^{2}}$ denotes $\chi^{2}$-capacity (Section 2.1).
The idea is that $C_{\chi^{2}}(P^{\star d})$ for a BMS channel $P$ is the expectation of a function involving $d$ i.i.d. variables. When $d$ is large, the result can be approximated using the expectation of a function involving a Gaussian variable. For the BOHT model, [GP23b] established the Gaussian approximation. The resulting function is a Gaussian integral (Eq. (24)), and we prove the numerical observation that it has no fixed points on $(0,1]$.
The second part (robust reconstruction) is more difficult. We show that for any $r\geq 2$, there exists $\epsilon=\epsilon(r)>0$ such that if $(r-1)d\lambda^{2}\leq 1$, then for any BMS channel $P$ with $C_{\chi^{2}}(P)\leq\epsilon$, we have
@@
There are several technical difficulties in the this approach. The first difficulty is that $C_{\chi^{2}}(\operatorname{\mathrm{BP}}(P))$ is not a function of $C_{\chi^{2}}(P)$. Any BMS $P$ is described by a random variable $\theta_{P}$ on $[-1,1]$ (see Section 2.1). The $\chi^{2}$-information $C_{\chi^{2}}(P)$ satisfies
@@
Because $\|\theta_{P}\|\leq 1$ almost surely for any $P$, we have $y_{k}\leq x_{k}$. We first perform a crude expansion of $x_{k+1}$ and $y_{k+1}$, which roughly speaking allows us to conclude that $y_{k}=O(x_{k}^{2})$. In this way we can control the influence of $y_{k}$ and perform a finer expansion of $x_{k+1}$, which give us Eq. (16).
The second difficulty is that, unlike previous works which work with trees, we work with hypertrees. As we noted, $C_{\chi^{2}}(\operatorname{\mathrm{BP}}(P))$ is the expectation of a function with $(r-1)d$ variables, while previous works have $d$ variables. When performing the expansion, there will be a large number of terms, even after considering symmetry. Our solution is to perform the expansion in two steps. Let $M^{\prime}_{k+1}$ be the channel $M_{k}^{\times(r-1)}\circ B_{r,\lambda}$. We decompose the BP recursion $M_{k+1}=\operatorname{\mathrm{BP}}(M_{k})$ into two conceptually smaller steps $M_{k}\mapsto M^{\prime}_{k+1}$ and $M^{\prime}_{k+1}\mapsto M_{k+1}$. Equivalently, we decompose the expansion step $(x_{k},y_{k})\mapsto(x_{k+1},y_{k+1})$ into two steps $(x_{k},y_{k})\mapsto(x^{\prime}_{k+1},y^{\prime}_{k+1})$ and $(x^{\prime}_{k+1},y^{\prime}_{k+1})\mapsto(x_{k+1},y_{k+1})$, where
@@
The fine expansions we get for $x$ and $x^{\prime}$ are as follows.